# Supernova Electron-Neutrino Interactions with Xenon in the nEXO Detector

### v1

@@ Abstract
Electron-neutrino charged-current interactions with xenon nuclei were modeled in the nEXO neutrinoless double-beta decay detector ($\sim$5-tonne, 90% 136Xe, 10% 134Xe) to evaluate its sensitivity to supernova neutrinos. Predictions for event rates and detectable signatures were modeled using the MARLEY event generator. We find good agreement between MARLEY’s predictions and existing theoretical calculations of the inclusive cross sections at supernova neutrino energies. The interactions modeled by MARLEY were simulated within the nEXO simulation framework and were run through an example reconstruction algorithm to determine the detector’s efficiency for reconstructing these events. The simulated data, incorporating the detector response, were used to study the ability of nEXO to reconstruct the incident electron-neutrino spectrum and these results were extended to a larger xenon detector of the same isotope enrichment. We estimate that nEXO will be able to observe electron-neutrino interactions with xenon from supernovae as far as 5 to 8 kpc from earth, while the ability to reconstruct incident electron-neutrino spectrum parameters from observed interactions in nEXO is limited to closer supernovae.
@@ I Introduction
The predicted rate of core-collapse supernovae (CCSN) within our galaxy is estimated to be a few per century, with large uncertainties [1, 2, 3, 4, 5]. When a CCSN occurs, $\sim$99of its gravitational binding energy is emitted in the form of neutrinos [6], which can both provide an early alert to astronomers that a supernova has occurred [7], as well as provide valuable information about the explosion dynamics. Neutrinos and antineutrinos of all flavors are expected to be produced during a core-collapse supernova, with the energy and flavor content varying throughout the explosion; in particular, electron neutrinos are expected to be produced in abundance during the infall phase and neutronization burst of a CCSN [8]. By studying the energy spectra of the different flavors of neutrinos produced during a CCSN, it is possible to image the interior dynamics of the collapsing star, as different flavors will decouple from thermal equilibrium at different depths. Studying the time, energy, and flavor composition of detected neutrinos can provide insight onto the neutrino mass ordering [9], set bounds on the neutrino mass [10], and test non-standard neutrino interactions and physics beyond the Standard Model [11, 12].
@@
Two examples of detectors with electron-neutrino charged-current ($\nu_{e}$) sensitivity are HALO [13] and the future DUNE experiment [14], which will utilize $\nu_{e}$ events on lead and argon, respectively, to search for supernovae. As $\nu_{e}$ cross sections tend to scale with increasing neutron excess in the interacting nucleus [15], smaller detectors with neutron-rich targets such as xenon can still have sensitivity to supernova electron neutrinos. By contrast, CC cross sections for electron antineutrinos are suppressed for heavier targets as a result of Pauli blocking [16]. While current xenon-based detectors have limited sensitivity to electron neutrinos from CCSN owing to their small masses, future detectors have been proposed for neutrinoless double beta-decay ($0\nu\beta\beta$) and WIMP dark-matter searches with masses ranging from 30-tonnes to 3-kilotonnes [17, 18, 19], that would have increased sensitivity.
The nEXO detector will search for neutrinoless double beta-decay using a single-phase liquid xenon time-projection chamber (TPC) enriched to 90% in the 136Xe isotope [20]. nEXO’s sensitivity to $0\nu\beta\beta$ decay is robust against backgrounds and “unknown unknowns” due to the use of multiple observables for signal-to-background discrimination. This includes, but is not limited to, its design energy resolution of $<$1% at the large 136Xe $Q$-value of 2.458 MeV. These factors result in an estimated $0\nu\beta\beta$ half-life sensitivity of $1.35\times 10^{28}$ years at 90% confidence level after 10 years of data taking. Many of the properties that make nEXO ideal for detecting $0\nu\beta\beta$ also give it sensitivity to supernova electron-neutrino CC interactions with xenon ($\nu_{e}$-Xe) such as its large homogeneous xenon volume, location deep underground, low-background design, optimization for MeV-scale physics, and planned 10-year exposure. Existing calculations [21, 22, 23] of $\nu_{e}$-Xe interactions typically focus on interaction rates, but not on how many of those interactions would be detectable. To realistically study these interactions in nEXO, particles produced by electron-neutrino interaction must be modeled along with the detector response to their interactions.
Sec. II describes the nEXO detector and relevant details for detecting supernovae. Sec. III focuses on the modifications of the MARLEY (Model of Argon Reaction Low Energy Yields) event generator [24] required for modeling $\nu_{e}$-Xe interactions; here MARLEY’s cross section predictions are compared with existing theoretical calculations. Sec. IV details the simulation of events predicted by MARLEY and the reconstruction algorithm employed to predict the visible spectrum nEXO will observe as a result of electron-neutrino interactions. Finally, Sec. V studies the ability of the nEXO detector to reconstruct the incident supernova electron-neutrino parameters. These results are extended to a larger 300-tonne detector of similar isotopic enrichment.
@@ II The nEXO detector
The conceptual design for the nEXO detector features $\sim$5-tonnes of xenon enriched to 90% in the isotope 136Xe. The remaining 10% consists mainly of 134Xe [20]. The liquid xenon will be contained within a cylindrical copper vessel with an inner diameter of 127.7 cm and height of 127.7 cm. Additional details on the detector can be found in Ref. [20]. The collaboration is planning to locate the detector deep underground at the SNOLAB Cryopit, with $\sim$6000 meters water-equivalent overburden, where the cosmic muon rate is 0.27 muons $\mbox{m}^{-2}\mbox{ day}^{-1}$ [25]. The detector will be surrounded by a 12.3 m-diameter 13.3 m-high water tank, acting as an active muon veto as well as providing passive shielding from external backgrounds. In addition to being sensitive to supernova electron-neutrino interactions in the xenon target, the detector will also potentially be sensitive to supernova electron-antineutrino IBD interactions in the water tank [7].
@@
Particle interactions in the xenon produce vacuum ultraviolet scintillation light, which will be detected by silicon photomultipliers (SiPMs) lining the vertical walls of the detector, with the number of optical photons (OPs) serving as the measure of light intensity. Simultaneously, particle interactions will ionize xenon atoms, producing electrons, which are subsequently drifted to the detector anode with an applied electric field of 400 V/cm. Once this charge reaches the anode, it is read out by an array of 120 charge tiles (10 cm $\times$ 10 cm) consisting of perpendicular x- and y- strips with a pitch of 6 mm [26]. Using the fast timing information provided by the scintillation light, along with the location of hits on the charge tile, three-dimensional reconstruction of the incident particle interactions is possible.
@@
While the supernova neutrino detection trigger configuration is still under development, this study uses a conservative scintillation light threshold based on the observed number of optical photons and the simulated light collection efficiency (see Ref. [20] for details) that corresponds to $\sim$100efficiency for 500 keV depositions. See Sec. IV.1 for more details on the example reconstruction algorithm used in this analysis and the study’s trigger threshold. The large number of electron-antineutrino interactions in the active water shield could potentially be used to trigger the TPC data acquisition to search for lower energy depositions in the xenon [7].
@@
Steady-state backgrounds are not a major concern for supernova detection in nEXO, owing to the large energy of $\nu_{e}$-Xe events, low muon flux, and short duration of the CCSN burst. Potential sources of background are pile-up of lower-energy events, and muons that do not trigger the veto system. For this study, we assume these backgrounds are negligible. The contribution from them and their expected contribution over the $\sim$10-second CCSN burst could be studied with steady-state data collected during nEXO’s planned 10-year exposure. Additionally, the detector could potentially search for coincidences delayed de-excitations of metastable states of 136Cs [27, 28, 29] to identify CC events and distinguish them from background, although this may be difficult for interactions at supernova neutrino energies owing to the small metastable de-excitation energy relative to the prompt deposited energy.
@@ III Modeling xenon charged-current interactions with MARLEY
As the first step toward simulating electron neutrino CC interactions with xenon nuclei in nEXO, the particles produced from these interactions are modeled using the MARLEY event generator [30]. These reactions as described by the following equation,
@@
where the asterisk indicates that the resulting cesium nucleus is typically expected to be generated in an excited state, which can subsequently de-excite through the emission of gamma rays, neutrons, and other particles. MARLEY models CC interactions in the allowed approximation (four-momentum transfer $q\rightarrow 0$, Fermi motion neglected); details on the underlying physics models used by MARLEY can be found in Ref. [24].
While originally designed for CC interactions on argon [31], MARLEY has been adapted for a variety of different nuclear targets [32, 33]. The required inputs for MARLEY are the Gamow-Teller (GT) and Fermi (F) strength distributions. The former can be measured in charge-exchange reactions such as $(p,n)$ and $({}^{3}\mbox{He},t)$, or can be calculated with a variety of theoretical approaches.
@@ III.1 Gamow-Teller and Fermi strengths
The experimentally-measured GT matrix elements are multiplied by a factor of $g_{A}^{2}=1.26^{2}$ to form the weak matrix elements, adopting the MARLEY input format, with the value of $g_{A}$ determined from the normalization assumed in Ref. [38]. While recent ab-initio calculations of matrix elements are able to reproduce $\beta$-decay rates [39], calculations with approximate nuclear models, such as those used within our study, tend to over-predict the GT strength [40]. We multiply these theoretical matrix elements by a quenching factor corresponding to $g_{A,\mbox{eff}}^{2}=(0.7)^{2}$. The quenched value of $g_{A}$ for these calculated matrix elements is similar to what is used in Ref. [35], and was chosen to compare MARLEY’s predictions to the calculations of Ref. [22]. Additionally, this value is in agreement with studies of $g_{A}$ quenching for targets in a similar mass range [41].
The Fermi strengths assume a value of $B(F)=g_{V}^{2}(N-Z)$ where $N$ is the number of neutrons in the target nucleus, $Z$ is the number of protons, and $g_{V}=1$ is the vector coupling constant. This Fermi strength is assumed to be located entirely at the isobaric analog state (IAS) of the product nucleus, which is typical for most nuclei with $N\gg{}Z$ [34]. For 136Xe this IAS has been measured to occur at an excitation energy of 13.386 MeV [34] relative to the 136Cs ground state. For 134Xe, the formalism in Ref. [34], Eq. 2, is used to predict the location of this state, $E_{X}^{IAS}$, which is estimated to be accurate to within a few hundred keV [34].
@@
Here $M(Z,A)$ and $M(Z+1,A)$ refers to the atomic masses of the initial and final nucleus, $M(n)$ is the mass of the neutron and $M(H)$ the mass of a hydrogen atom. The calculated location of the isobaric analog state in 134Cs is 12.189 MeV, using the mass evaluation from Refs. [42, 43]. As a check of the validity of this equation in this mass range, we calculate the value of the IAS of 136Xe in 136Cs via Eq. 2 and obtain a value of 13.258 MeV, in good agreement with the measured value. The measured location is used in the MARLEY input file.
For this study, we focus on the isotopes of primary interest to the nEXO detector. This work could be extended to cover other stable naturally-occurring isotopes of xenon if experimental or theoretical Gamow-Teller strength predictions were readily available, particularly for stable non-even-even isotopes.
III.2 De-excitation data
MARLEY utilizes de-excitation data from TALYS-1.6 [44] to predict the observed discrete gammas emitted from a $\nu_{e}$-Xe interaction. While data exists for 134Cs excited states, until recently there was little experimental data on the excited states of 136Cs. Data from Ref. [28, 45] are added to the de-excitation data MARLEY uses; data from Ref. [34] are used for higher energy discrete states. We follow the approach from TALYS for spin-parity assignment when those quantities are unknown for a specific excited state [46]. To assign spin, a histogram is generated from the spins of lower-energy states and compared to a Wigner distribution, and spin is assigned based on the bin that most underestimates the Wigner distribution, using a spin cut-off parameter described by Eq. 237 within Ref. [46]. To assign parity, the distribution of parities of states below the current level is generated, and parity is assigned to balance that distribution. The MARLEY approach of using a Standard Lorentzian to model de-excitations is followed to generate predictions for branching ratios [24]. Improved measurements of the spin-parity of these excited states and their branching ratios would lead to more accurate predictions of experimental signatures in the future, although this is expected to have a small impact on the predicted energy distribution from $\nu_{e}$-Xe interactions in nEXO.
The MARLEY generator produces all output particles from a CC interaction simultaneously. Modifications are needed for a detailed study of the possibility of tagging the short-lived metastable states in 136Cs [28, 29] that can be produced as a result of the CC interaction.
III.3 Other electron-neutrino interactions
@@
$\nu_{e}+e^{-}\rightarrow\nu_{e}^{\prime}+e^{-\prime}$ (3)
can produce MeV-scale signals, but has a smaller cross section than CC scattering off xenon nuclei. We include neutrino-electron elastic scattering from electron neutrinos in the events we simulated. Contributions from electron antineutrinos, as well as other neutrino flavors through inelastic NC interactions, are not included. Electron-antineutrino CC interactions with xenon are expected to have cross sections that are approximately two orders of magnitude lower than those for electron neutrinos [22]. Inelastic NC cross sections are also expected to be smaller than those from electron-neutrino charged-current scattering [49], although the interaction rate is increased from the additional neutrino flavors that can interact through this channel. Modeling inelastic NC interactions with xenon in nEXO is beyond the scope of this paper, and could be a potential future area of study.

### v2

@@ Abstract
Electron-neutrino charged-current interactions with xenon nuclei were modeled in the nEXO neutrinoless double-$\beta$ decay detector ($\sim$5 metric ton, 90% 136Xe, 10% 134Xe) to evaluate its sensitivity to supernova neutrinos. Predictions for event rates and detectable signatures were modeled using the MARLEY (Model of Argon Reaction Low Energy Yields) event generator. We find good agreement between MARLEY’s predictions and existing theoretical calculations of the inclusive cross sections at supernova neutrino energies. The interactions modeled by MARLEY were simulated within the nEXO simulation framework and were run through an example reconstruction algorithm to determine the detector’s efficiency for reconstructing these events. The simulated data, incorporating the detector response, were used to study the ability of nEXO to reconstruct the incident electron-neutrino spectrum and these results were extended to a larger xenon detector of the same isotope enrichment. We estimate that nEXO will be able to observe electron-neutrino interactions with xenon from supernovae as far as 5-8 kpc from Earth, while the ability to reconstruct incident electron-neutrino spectrum parameters from observed interactions in nEXO is limited to closer supernovae.
@@ I Introduction
The predicted rate of core-collapse supernovae (CCSNe) within our Galaxy is estimated to be a few per century, with large uncertainties [1, 2, 3, 4, 5]. When a CCSN occurs, $\sim$99of its gravitational binding energy is emitted in the form of neutrinos [6], which can both provide an early alert to astronomers that a supernova has occurred [7], as well as provide valuable information about the explosion dynamics. Neutrinos and antineutrinos of all flavors are expected to be produced during a core-collapse supernova, with the energy and flavor content varying throughout the explosion. In particular, electron neutrinos are expected to be produced in abundance during the infall phase and neutronization burst of a CCSN [8]. By studying the energy spectra of the different flavors of neutrinos produced during a CCSN, it is possible to image the interior dynamics of the collapsing star, as different flavors will decouple from thermal equilibrium at different depths. Studying the time, energy, and flavor composition of detected neutrinos can provide insight into the neutrino mass ordering [9], set bounds on the neutrino mass [10], and test nonstandard neutrino interactions and physics beyond the Standard Model [11, 12].
@@
Two examples of detectors with electron-neutrino charged-current ($\nu_{e}$) sensitivity are HALO [13] and the future DUNE experiment [14], which will utilize $\nu_{e}$ events on lead and argon, respectively, to search for supernovae. As $\nu_{e}$ cross sections tend to scale with increasing neutron excess in the interacting nucleus [15], smaller detectors with neutron-rich targets such as xenon can still have sensitivity to supernova electron neutrinos. Additionally, future detectors have been proposed for neutrinoless double-$\beta$ decay ($0\nu\beta\beta$) and weakly interacting massive particle dark-matter searches with masses ranging from 30 metric tons to 3 metric kt [16, 17, 18], which would have increased sensitivity.
The nEXO detector will search for neutrinoless double-$\beta$ decay using a single-phase liquid xenon time-projection chamber (TPC) enriched to 90% in the 136Xe isotope [19]. nEXO’s sensitivity to $0\nu\beta\beta$ decay is robust against backgrounds and “unknown unknowns” due to the use of multiple observables for signal-to-background discrimination. This includes, but is not limited to, its design energy resolution of $<$1% at the large 136Xe $Q$ value of 2.458 MeV. These factors result in an estimated $0\nu\beta\beta$ half-life sensitivity of $1.35\times 10^{28}$ yr at 90% confidence level after 10 yr of data taking. Many of the properties that make nEXO ideal for detecting $0\nu\beta\beta$ also give it sensitivity to supernova electron-neutrino CC interactions with xenon ($\nu_{e}$-Xe) such as its large homogeneous xenon volume, location deep underground, low-background design, optimization for MeV-scale physics, and planned 10-yr exposure. Existing calculations [20, 21, 22] of $\nu_{e}$-Xe interactions typically focus on interaction rates, but not on how many of those interactions would be detectable. To realistically study these interactions in nEXO, particles produced by electron-neutrino interaction must be modeled along with the detector response to their interactions.
Section II describes the nEXO detector and relevant details for detecting supernovae. Section  III focuses on the modifications of the MARLEY (Model of Argon Reaction Low Energy Yields) event generator [23] required for modeling $\nu_{e}$-Xe interactions; here MARLEY’s cross section predictions are compared with existing theoretical calculations. Section IV details the simulation of events predicted by MARLEY and the reconstruction algorithm employed to predict the visible spectrum nEXO will observe as a result of electron-neutrino interactions. Finally, Sec. V studies the ability of the nEXO detector to reconstruct the incident supernova electron-neutrino parameters. These results are extended to a larger 300 metric ton detector of similar isotopic enrichment.
@@ II The nEXO detector
The conceptual design for the nEXO detector features $\sim$5 metric tons of xenon enriched to 90% in the isotope 136Xe. The remaining 10% consists mainly of 134Xe [19]. The liquid xenon will be contained within a cylindrical copper vessel with an inner diameter of 127.7 cm and height of 127.7 cm. Additional details on the detector can be found in Ref. [19]. The collaboration is planning to locate the detector deep underground at the SNOLAB Cryopit, with $\sim$6,000-m water-equivalent overburden, where the cosmic muon rate is 0.27 muons $\mbox{m}^{-2}\mbox{ day}^{-1}$ [24]. The detector will be surrounded by a 12.3-m-diameter 13.3-m-high water tank, acting as an active muon veto as well as providing passive shielding from external backgrounds. In addition to being sensitive to supernova electron-neutrino interactions in the xenon target, the detector will also potentially be sensitive to supernova electron-antineutrino IBD interactions in the water tank [7].
@@
Particle interactions in the xenon produce vacuum ultraviolet scintillation light, which will be detected by silicon photomultipliers (SiPMs) lining the vertical walls of the detector, with the number of optical photons (OPs) serving as the measure of light intensity. Simultaneously, particle interactions will ionize xenon atoms, producing electrons, which are subsequently drifted to the detector anode with an applied electric field of 400 V/cm. Once this charge reaches the anode, it is read out by an array of 120 charge tiles (10 $\times$ 10 cm) consisting of perpendicular x and y strips with a pitch of 6 mm [25]. Using the fast timing information provided by the scintillation light, along with the location of hits on the charge tile, three-dimensional reconstruction of the incident particle interactions is possible.
@@
While the supernova neutrino detection trigger configuration is still under development, this study uses a conservative scintillation light threshold based on the number of collected OPs in the xenon volume. The collected number of OPs are a function of the number of produced OPs and the light collection efficiency (see Ref. [19] for details). Our chosen threshold corresponds to $\sim$100efficiency for 500 keV depositions. See Sec. IV.1 for more details on the example reconstruction algorithm used in this analysis and this study’s trigger threshold. The large number of expected electron-antineutrino interactions in the active water shield could also be used to trigger the TPC data acquisition to search for lower-energy depositions in the xenon [7].
@@
Steady-state backgrounds are not a major concern for supernova detection in nEXO, owing to the large energy of $\nu_{e}$-Xe events, low muon flux, and short duration of the CCSN burst. Potential sources of background are pileup of lower-energy events, and muons that do not trigger the veto system. For this study, we assume these backgrounds are negligible. The contribution from them and their expected contribution over the $\sim$10-sec CCSN burst could be studied with steady-state data collected during nEXO’s planned 10-yr exposure.
@@ III Modeling xenon charged-current interactions with MARLEY
As the first step toward simulating electron-neutrino CC interactions with xenon nuclei in nEXO, the particles produced from these interactions are modeled using the MARLEY event generator [26]. These reactions as described by the following equation,
@@
where the asterisk indicates that the resulting cesium nucleus is typically expected to be generated in an excited state, which can subsequently deexcite through the emission of $\gamma$ rays, neutrons, and other particles. MARLEY models CC interactions in the allowed approximation (four-momentum transfer $q\rightarrow 0$, Fermi motion neglected); details on the underlying physics models used by MARLEY can be found in Ref. [23].
While originally designed for CC interactions on argon [27], MARLEY has been adapted for a variety of different nuclear targets [28, 29]. The required inputs for MARLEY are the Gamow-Teller (GT) and Fermi (F) strength distributions. The former can be measured in charge-exchange reactions, such as $(p,n)$ and $({}^{3}\mbox{He},t)$, or can be calculated with a variety of theoretical approaches.
@@ III.1 Gamow-Teller and Fermi strengths
The experimentally measured GT matrix elements are multiplied by a factor of $g_{A}^{2}=1.26^{2}$ to form the weak matrix elements, adopting the MARLEY input format, with the value of $g_{A}$ determined from the normalization assumed in Ref. [34]. While recent ab initio calculations of matrix elements are able to reproduce $\beta$-decay rates [35], calculations with approximate nuclear models, such as those used within our study, tend to overpredict the GT strength [36]. We multiply these theoretical matrix elements by a quenching factor corresponding to $g_{A,\mbox{eff}}^{2}=(0.7)^{2}$. The quenched value of $g_{A}$ for these calculated matrix elements is similar to what is used in Ref. [31], and was chosen to compare MARLEY’s predictions to the calculations of Ref. [21]. Additionally, this value is in agreement with studies of $g_{A}$ quenching for targets in a similar mass range [37].
The Fermi strengths assume a value of $B(F)=g_{V}^{2}(N-Z)$ where $N$ is the number of neutrons in the target nucleus, $Z$ is the number of protons, and $g_{V}=1$ is the vector coupling constant. This Fermi strength is assumed to be located entirely at the isobaric analog state (IAS) of the product nucleus, which is typical for most nuclei with $N\gg{}Z$ [30]. For 136Xe this IAS has been measured to occur at an excitation energy of 13.386 MeV [30] relative to the 136Cs ground state. For 134Xe, the formalism in Ref. [30], Eq. 2, is used to predict the location of this state, $E_{X}^{IAS}$, which is estimated to be accurate to within a few hundred keV [30],
@@
Here $M(Z,A)$ and $M(Z+1,A)$ refer to the atomic masses of the initial and final nucleus, $M(n)$ is the mass of the neutron, and $M(H)$ is the mass of a hydrogen atom. The calculated location of the isobaric analog state in 134Cs is 12.189 MeV, using the mass evaluation from Refs. [38, 39]. As a check of the validity of this equation in this mass range, we calculate the value of the IAS of 136Xe in 136Cs via Eq. 2 and obtain a value of 13.258 MeV, in good agreement with the measured value. The measured location is used in the MARLEY input file.
For this study, we focus on the isotopes of primary interest to the nEXO detector. This work could be extended to cover other stable naturally occurring isotopes of xenon if experimental or theoretical Gamow-Teller strength predictions were readily available, particularly for stable non even-even isotopes.
III.2 Deexcitation data
MARLEY utilizes deexcitation data from TALYS 1.6 [40] to predict the observed discrete $\gamma$’s emitted from a $\nu_{e}$-Xe interaction. While data exist for 134Cs excited states, until recently there were little experimental data on the excited states of 136Cs. Data from Ref. [41, 42] are added to the deexcitation data MARLEY uses; data from Ref. [30] are used for higher-energy discrete states. We follow the approach from TALYS for spin-parity assignment when those quantities are unknown for a specific excited state [43]. To assign spin, a histogram is generated from the spins of lower-energy states and compared to a Wigner distribution, and spin is assigned based on the bin that most underestimates the Wigner distribution, using a spin cutoff parameter described by Eq. 237 within Ref. [43]. To assign parity, the distribution of parities of states below the current level is generated, and parity is assigned to balance that distribution. The MARLEY approach of using a standard Lorentzian to model deexcitations is followed to generate predictions for branching ratios [23]. Improved measurements of the spin-parity of these excited states and their branching ratios would lead to more accurate predictions of experimental signatures in the future, although this is expected to have a small impact on the predicted energy distribution from $\nu_{e}$-Xe interactions in nEXO.
III.3 Other neutrino interactions in nEXO
@@
[…]
Inelastic NC neutrino-nucleus scattering can produce MeV-scale $\gamma$’s in nEXO and is sensitive to all flavors of supernova neutrinos. However, this process is predicted to have a smaller cross section than that of $\nu_{e}$-Xe [46]. Modeling these types of interactions in nEXO is beyond the scope of this paper, and could be a potential future area of study.
$\nu eES$, described by Eq. 3,
$\nu+e^{-}\rightarrow\nu^{\prime}+e^{-\prime},$ (3)
can produce MeV-scale signals, but is expected to have a smaller cross section than that of $\nu_{e}$-Xe. The total cross section for this process depends on the flavor of the interacting neutrino, with the largest corresponding to $\nu_{e}$s. We have included $\nu_{e}-e^{-}$ elastic scattering in our simulations, but have not included contributions from $\bar{\nu}_{e}$s and $\nu_{x}$. The $\nu eES$ cross section for $\bar{\nu}_{e}$ is smaller by a factor of $\sim$2, and for $\nu_{x}$ it is smaller by a factor of $\sim$4.
Electron antineutrinos can also interact through charged-current scattering on xenon, although the cross section for this interaction is expected to be approximately 2 orders of magnitude lower than that of electron neutrinos [21], as a result of Pauli blocking [47].
Finally, as discussed in Ref. [48], it may be possible to tag $\nu_{e}$-Xe interactions through timing coincidences with short-lived metastable states in the resulting cesium nucleus, such as those recently identified for 136Cs [41, 49]. This would enable separating $\nu_{e}$-Xe events from other types of neutrino interactions. This is not incorporated into our simulation, and may be difficult at supernova neutrino energies owing to the small metastable deexcitation energy relative to the prompt energy deposition.